A multiple regression model involves 6 independent variables and a sample of 20 data points. If we want to test the validity of the entire model at the 5% significance level, the critical value is: A) 2.92 B) 2.90 C) 2.85 D) 3.06
Answer #1

We apply F statistic in this case

Numerator degrees of freedom (k) = # of independent variables = 6

Denominator degrees of freedom = n - k - 1 = 20 - 6 - 1 = 13

At 5% sinificance level, alpha = 0.05

F statistic critical value = 2.915 = 2.92

So, option A) is correct